Lesson 12. Questions 1 to 5—convert the following number of moles to number of particles:

  1. 2.0 moles of CaCl2
  2. 1.5 moles of copper(II) ions
  3. 0.01 moles of H2O
  4. 5.0 moles of gold atoms
  5. 0.0025 moles of HCl
  6. What is the mass of 1 mole of Mg atoms?
  7. What is the molar mass of aluminum?
  8. What is the mass of 1 mole of oxygen atoms?
  9. What is the mass of 1 mole of oxygen molecules, O2?
  10. Calculate the molar mass of NaOH.
  11. What is the mass of 1.0 mole of CaCl2?
  12. Calculate the molar mass of Cu(NO3)2.
  13. Calculate the molar mass of Al2(SO4)3.
1
Expert's answer
2021-03-22T06:27:24-0400

Solution:

One mole of any substance contains 6.022×1023 particles.

Hence,

(1) 2.0 moles of CaCl2:

Number of CaCl2 molecules = (2.0 mol) × (6.022×1023 molecules / 1 mol) = 1.20×1024 CaCl2 molecules

Number of CaCl2 molecules = 1.20×1024 molecules


(2) 1.5 moles of copper(II) ions:

Number of copper(II) ions = (1.5 mol) × (6.022×1023 ions / 1 mol) = 9.03×1023 Cu2+ ions

Number of copper(II) ions = 9.03×1023 ions


(3) 0.01 moles of H2O:

Number of H2O molecules = (0.01 mol) × (6.022×1023 molecules / 1 mol) = 6.02×1021 H2O molecules

Number of H2O molecules = 6.02×1021 molecules


(4) 5.0 moles of gold atoms:

Number of gold atoms = (5.0 mol) × (6.022×1023 atoms / 1 mol) = 3.01×1024 Au atoms

Number of gold atoms = 3.01×1024 atoms


(5) 0.0025 moles of HCl:

Number of HCl molecules = (0.0025 mol) × (6.022×1023 molecules / 1 mol) = 1.506×1021 HCl molecules

Number of HCl molecules = 1.51×1021 molecules



(6) What is the mass of 1 mole of Mg atoms?

One mole of magnesium (Mg) atoms has a mass of 24.31 g.


(7) What is the molar mass of aluminum?

The molar mass of aluminum (Al) is 26.982 g/mol.


(8) What is the mass of 1 mole of oxygen atoms?

One mole of oxygen (O) atoms has a mass of 16 g.


(9) What is the mass of 1 mole of oxygen molecules, O2?

One mole of oxygen molecules (O2) has a mass of 32 g.


(10) Calculate the molar mass of NaOH.

M(Na) = 22.989 g/mol

M(O) = 15.999 g/mol

M(H) =1.0079 g/mol

So, its molar mass is 22.989 + 15.999 + 1.0079 = 39.996 g/mol

M(NaOH) = 39.996 g/mol


(11) What is the mass of 1.0 mole of CaCl2?

The molar mass of CaCl2 is 110.98 g/mol.

So, 1.0 mole of CaCl2 has a mass of 110.98 g.


(12) Calculate the molar mass of Cu(NO3)2.

M(Cu) = 63.546 g/mol

M(N) = 14.0067 g/mol

M(O) = 15.999 g/mol

So, its molar mass is 63.546 + 2×14.0067 + 6×15.999 = 187.5534 g/mol = 187.55 g/mol

M(Cu(NO3)2) = 187.55 g/mol


(13) Calculate the molar mass of Al2(SO4)3.

M(Al) = 26.98154 g/mol

M(S) = 32.065 g/mol

M(O) = 15.999 g/mol

So, its molar mass is 2×26.98154 + 3×32.065 + 12×15.999 = 342.146 g/mol = 342.15 g/mol

M(Al2(SO4)3) = 342.15 g/mol